jhonyy9 (jhonyy9):

prove by math induction : 1^2 +2^2 +3^2 + ... + n^2 = [n(n+1)(2n+1)]/6

jhonyy9 (jhonyy9):

first step n=1 1^2 = [1(1+1)(2*1+1)]/6 1 = [1*2*3]/6 1 = 6/6 1 = 1 so this is true rewrite it for n = k and will get 1^2 +2^2 + ...+k^2 = [k(k+1)(2k+1)]/6 suppose this is true and rewrite it for k=k+1 and prove it that is true -

OpenStudy (s4sensitiveandshy):

\[1^2 +2^2 +3^2 + ... + \color{red}{ n^2} = \frac{\color{Red}{ n}(\color{Red}{n}+1)(2\color{Red}{n}+1) }{ 6 }\] \[1^2 +2^2 +3^2 + ... + \color{red}{ k^2} = \frac{\color{Red}{ k}(\color{Red}{k}+1)(2\color{Red}{k}+1) }{ 6 }\] n=k+1 \[1^2 +2^2 +3^2 + ... + \color{red}{ k^2}+ \color{blue}{(k+1)^2} = \frac{\color{blue}{ (k+1)}[\color{Red}{(k+1)}+1][2\color{Red}{(k+1)}+1] }{ 6 }\]

OpenStudy (s4sensitiveandshy):

\[ \color{green}{1^2 +2^2 +3^2 + ... + \color{red}{ k^2}} +\color{blue}{(k+1)^2} = \frac{\color{blue}{( k+1)}[\color{Red}{(k+1)}+1][2\color{Red}{(k+1)}+1] }{ 6 }\] 1^2+2^2+3^2+......+k^2 = \[ 1^2+2^2+3^2+......+k^2 = \frac{ k(k+1)(2k+1) }{ 6 }\] thus.... \[ \color{orange}{ \frac{ k(k+1)(2k+1) }{ 6 }} +\color{blue}{(k+1)^2} = \frac{\color{blue}{( k+1)}[\color{Red}{(k+1)}+1][2\color{Red}{(k+1)}+1] }{ 6 }\]

OpenStudy (sshayer):

adding both sides (k+1)^2 \[1^2+2^2+...+k^2+(k+1)^2=\frac{ k(k+1)(2k+1 }{ 6 }+(k+1)^2\] \[=(k+1)\left[ \frac{ k(2k+1)+6(k+1) }{ 6 } \right]=(k+1)\left[ \frac{ 2k^2+k+6k+6 }{ 6 } \right]\] \[=(k+1)\left[ \frac{ 2k^2+7k+6 }{ 6 } \right]=(k+1)\left[\frac{ 2k^2+4k+3k+6 }{ 6 } \right]\] \[=(k+1)\left[ \frac{ 2k(k+2)+3(k+2) }{ 6 } \right]=\frac{ (k+1)(k+2)(2k+3) }{ 6 }\] \[=\frac{ (k+1)\left\{ \left( k+1 \right)+1 \right\}\left\{ 2(k+1)+1 \right\} }{ 6 }\]

OpenStudy (mww):

People often forget to use LHS and RHS which is important when you are comparing the given induction to the simulated induction (i.e. using the n = k result). Should technically add '= RHS' for completion.

jhonyy9 (jhonyy9):

@ParthKohli your opinion please about this above wrote proof by math induction is acceptably ? thank you

Parth (parthkohli):

Yes, it's completely correct.